# Gravity from (A)dS Yang-Mills: a natural reduction

### v2

@@ 4.2 Recovering gravitational dynamics
We are now able to show that the Euler-Lagrange equations of the Yang-Mills action in (2.8), reduced to  (3.18), (3.19), are consistent with gravitational equations once the gauge fields are identified as described in (4.13).
Equations (3.18)-(3.19), in components, read:
@@
$\left(\frac{1}{2}\mathcal{D}_{\sigma}^{\varpi}\tilde{\mathfrak{T}}^{a}_{\rho\lambda}+\frac{1}{4}\mathfrak{R}_{[\mu\nu]}^{ab}\vartheta_{b\sigma}\varepsilon^{\mu\nu}_{\rho\lambda}+\frac{\lambda}{2}\vartheta_{b\sigma}\vartheta^{a}_{\mu}\vartheta^{b}_{\nu}\varepsilon^{\mu\nu}_{\rho\lambda}\right)\varepsilon^{\alpha\sigma\rho\lambda}$ $=$ $\chi\Theta^{a}_{[\sigma\rho\lambda]}\varepsilon^{\alpha\sigma\rho\lambda}$ (4.15)
$\frac{1}{4}\left(\tilde{\mathfrak{T}}^{b}_{\rho\lambda}\vartheta^{a}_{\sigma}-\tilde{\mathfrak{T}}^{a}_{\rho\lambda}\vartheta^{b}_{\sigma}\right)\varepsilon^{\alpha\sigma\rho\lambda}$ $=$ $\chi\Sigma^{ab}_{[\sigma\rho\lambda]}\varepsilon^{\alpha\sigma\rho\lambda}$ (4.16)
@@
where we defined $\tilde{\mathfrak{T}}^{a}_{\rho\lambda}=\varepsilon^{\mu\nu}_{\rho\lambda}\mathfrak{T}^{a}_{[\mu\nu]}$ and we used the explicit expressions [$\mathfrak{R}^{ab}_{[\mu\nu]}=\partial_{\mu}\varpi^{ab}_{\nu}-\partial_{\nu}\varpi^{ab}_{\mu}+\varpi^{a}_{\mu\,c}\varpi^{cb}_{\nu}-\varpi^{a}_{\nu\,c}\varpi^{cb}_{\mu}$, $\mathfrak{T}^{a}_{[\mu\nu]}=\partial_{\mu}\vartheta^{a}_{\nu}-\partial_{\nu}\vartheta^{a}_{\mu}+\varpi^{a}_{\mu\,c}\vartheta^{c}_{\nu}-\varpi^{a}_{\nu\,c}\vartheta^{c}_{\mu}$.
Now, using the definitions of the sources given in (3.6) and (3.7), when the fields acquire a geometric interpretation Eqs. (4.15),(4.16) can be written as
$R^{a}_{\beta}-\frac{1}{4}\theta^{a}_{\beta}R$ $=$ ${\chi}(\mathscr{T}^{a}_{\beta}-\frac{1}{4}\theta^{a}_{\beta}\mathscr{T})+{\chi}(s^{a}_{\beta}-\frac{1}{4}\theta^{a}_{\beta}s)$ (4.17)
$\tilde{T}^{a}_{\lambda\tau}$ $=$ $-{\chi}\left(\mathfrak{S}^{a}_{\lambda\tau}+\frac{1}{2}\mathfrak{S}_{\lambda}\theta^{a}_{\tau}-\frac{1}{2}\mathfrak{S}_{\tau}\theta^{a}_{\lambda}\right)$ (4.18)
where $\lambda$ in (4.15) is replaced by the trace of the same equation. Here we have introduced
@@
$s^{a}_{\beta}=-\frac{1}{2}\varepsilon^{\sigma\rho\lambda}_{\beta}\mathcal{D}^{\omega}_{\sigma}\left(\mathfrak{S}^{a}_{\rho\lambda}+\frac{1}{2}\mathfrak{S}_{\rho}\theta^{a}_{\lambda}-\frac{1}{2}\mathfrak{S}_{\lambda}\theta^{a}_{\rho}\right)$ (4.19)
with $s=\theta^{\beta}_{a}{s}^{a}_{\beta}$ its trace, given $\mathfrak{S}^{a}_{\lambda\tau}=\mathfrak{S}^{\alpha}_{cd}\theta^{c}_{\lambda}\theta^{d}_{\tau}\theta^{a}_{\alpha}$ with $\mathfrak{S}^{\alpha}_{cd}=\varepsilon_{cdab}\,\mathcal{S}^{ab\,\alpha}$ and $\mathfrak{S}_{\lambda}=\mathfrak{S}^{a}_{\lambda\tau}\theta^{\tau}_{a}$ its partial trace. Moreover, $R^{a}_{\beta}=\theta^{\nu}_{b}R^{ab}_{[\beta\nu]}$ is the Ricci tensor, $R=\theta^{\beta}_{a}\theta^{\nu}_{b}R^{ab}_{[\beta\nu]}$ is the scalar curvature, while $\mathscr{T}=\theta^{\beta}_{a}\mathscr{T}^{a}_{\beta}$ is the trace of the energy-momentum tensor.
This shows that in the general case where the torsion is algebraically related with spin, the equations of motion can be associated with an extended unimodular gravity framework, characterized by an additional spin contribution to the energy-momentum tensor. Whereas, in the case of zero spin and zero torsion, the equations are exactly those of unimodular gravity [31]. Indeed, in this case, we get
$R^{a}_{\beta}-\frac{1}{4}\theta^{a}_{\beta}R$ $=$ ${\chi}(\mathscr{T}^{a}_{\beta}-\frac{1}{4}\theta^{a}_{\beta}\mathscr{T})$ (4.20)
$\tilde{T}^{a}_{\lambda\tau}$ $=$ 00 0 (4.21)
A further contraction with $\eta_{ab}\theta^{b}_{\mu}$ allows to express equation (4.20) in terms of the induced metric field $g_{\beta\mu}=\eta_{ab}\theta^{a}_{\beta}\theta^{b}_{\mu}$ as
$R_{\beta\mu}-\frac{1}{4}g_{\beta\mu}R={\chi}(\mathscr{T}_{\beta\mu}-\frac{1}{4}g_{\beta\mu}\mathscr{T}).$ (4.22)
@@ 5 Discussion
The proposed model recovers gravitational dynamics as a specific sector of a (A)dS Yang-Mills theory while reducing the gauge fields to geometric tetrad and spin connection via Inönü-Wigner contraction of the (A)dS algebra to the Poincaré algebra, without modifying the action functional. In this framework, the symmetries of the gauge theory map to transformations of the geometric fields of first-order gravity, while a specific sector of the gauge dynamics, identified by an algebraic relationship between torsion and spin, corresponds to the dynamics of a generalized unimodular gravity theory (to be compared e.g. with [32]). The non-propagating torsion in the model is well motivated both theoretically and experimentally. For instance, this is a key feature of gauge theory approaches to gravity, such as Einstein-Cartan and Einstein-Sciama-Kibble theories [33].
As a first remark, it should be noted that our derivation of the equations differs from the standard approach of unimodular gravity[31], where the presence of a cosmological constant is not postulated, but derived as an integration constant. In our approach, the cosmological constant is present as an inherent scale of the theory. Indeed, as a consequence of the trace computation used in the passage from equation (4.15) to (4.17), in the case of $s=0$, the following equation is obtained
$\frac{R}{4}=\Lambda-\frac{\chi}{4}\mathscr{T}$ (5.1)
where $\Lambda=-3\lambda+\chi{{\mathscr{T}}}/{2}$. The expression above allows to explicitly introduce the cosmological constant in the equations of motion. Eq. (4.20) becomes then
$R^{a}_{\beta}-\frac{1}{2}\theta^{a}_{\beta}R+\theta^{a}_{\beta}\Lambda=\chi\mathscr{T}^{a}_{\beta}.$ (5.2)
A similar relation can be obtained in the case of spin nonzero.
@@
As for the parameters of the theory, in the limit $\alpha\rightarrow 0$ we have $\chi=-{2g^{2}}/{\hbar\lambda}={16\pi G_{N}}/{c^{3}}$. It follows that $\lambda$ and $g^{2}$ can be expressed in relation to one another through the squared Planck length $l^{2}_{p}=\hbar G_{N}/c^{3}$. As a consequence of the relation between $g^{2}$ and $\lambda$ and the convention adopted in the construction of the action, $\lambda$ must be negative, which in turn constrains $\alpha$ to be negative, as shown in Eq (2.2). This selection leads to $SO(2,3)$ as the family gauge structure group. However, adopting a different sign convention for the action, would have led to a positive $\alpha$, thereby selecting the $SO(1,4)$ family. This highlights the fact that the sign of $\alpha$ does not play an essential role in the construction, which can be developed in a fully general manner for both the de Sitter and anti-de Sitter cases.
@@
A few more comments are in order. As discussed in [31], the reliance of modern physics on two closely related but separate principles, i.e. general covariance and gauge invariance, may be unsatisfactory. Unifying these principles into a single framework is a task, and two main approaches emerge: one views gauge symmetry as a manifestation of general covariance in higher dimensions, while the other interprets general covariance as a form of gauge symmetry. The results presented in this paper go in the direction of the second approach. In fact the gauge transformations of $\vartheta$ generate diffeomorphisms. This occurs in the limit $\alpha\rightarrow 0$, where the gauge nature of the theory changes, providing a geometric interpretation of gauge potential components as the tetrad and the Lorentz connection.
@@
Since the model is obtained from a standard Yang-Mills action on Minkowski space-time, well established quantization techniques of non-Abelian gauge theories apply, although the issues related to the non-compactness of the structure group remain open, both in terms of unitarity and renormalizability of the quantum theory [21]. The results for the gravity sector of the theory should be analised while performing the limit $\alpha\rightarrow 0$.
@@
The possible physical meaning of the proposed mechanism, which reduces the gauge dynamics to the gravitational sector, deserves further investigation. One possibility could be a Higgs-like mechanism, which would involve a spontaneous breaking of the gauge symmetry to the residual Lorentz gauge symmetry, independently of the $\alpha$ limit (see e.g. [34, 35]). Another possibility concerns the idea that the gauge parameters of the theory, $\lambda$ and $g$, be renormalised during quantization and would exhibit a running dependence on $\alpha$. The latter, as a dimensionless parameter, could be linked to the energy scale only if the Planck length is considered as a fundamental scale. A running of the gauge parameters $\lambda$ and $g$ could cause them to weigh the various tensors in the Euler-Lagrange equations differently, making some of them negligible compared to others. In other words, the dynamics we have described could emerge as effective dynamics resulting from the quantum nature of the theory. Finally, the meaning of the contraction parameter $\alpha$ and its interplay with the cosmological constant have to be understood, also in relation with the renormalization group flow.
As a final speculation, we observe that, since the dynamical metric is quadratic in the sector of P-translations of the gauge algebra, the undelying gauge theory could be a natural candidate for a double copy formulation of gravity [36, 37, 38]. We plan to address this hypothesis in future work.
@@ Appendix A The (A)dS groups
In this appendix we fix the notation and our convention for the structure groups $SO({p},{q})$ of the Yang-Mills model analysed in the paper. These are the pseudo-orthogonal groups $SO({p},{q})$ with $p+q=5$, namely the special Lie groups of linear transformations of a real $5$-dimensional vector space that leave invariant a non-degenerate symmetric bilinear form $\eta$ of signature $(p,q)$. The algebra $\mathfrak{so}(p,q)$ is provided by generators $M_{AB}$, $A,B=0,\dots,4$, and Lie brackets
$[M_{AB},M_{CD}]=\eta_{AD}M_{BC}-\eta_{AC}M_{BD}+\eta_{BC}M_{AD}-\eta_{BD}M_{AC}:=f^{KL}_{AB,CD}M_{KL}$ (A.1)
@@
Specifically we are interested in the de Sitter group $SO(1,4)$ preserving the quadratic form $\eta_{AB}=diag(1,-1,-1,-1,-1)$ and the anti-de Sitter group $SO(2,3)$ preserving $\eta_{AB}=diag(1,-1,-1,-1,1)$. Our choice of basis for their Lie algebra is the one which singles out the Lorentz subalgebra. Indeed, if $a,b=0,1,2,3$ we can write
@@
where + and -- refer respectively to the de Sitter and anti-de Sitter algebra. Hence we define
@@
where $\eta_{ab}$ is the Minkowski metric in 4d with signature (+,,,)(+,-,-,-)( + , - , - , - ) while $K^{cd}_{a,b}=\frac{1}{2}(\delta^{c}_{a}\delta^{d}_{b}-\delta^{c}_{b}\delta^{d}_{a})$. The parameter $\alpha$ can take positive, negative, or zero values. The Poincaré algebra $\mathfrak{iso}(1,3)$ can be derived from $\mathfrak{so}(1,4)$ or $\mathfrak{so}(2,3)$ in the limit of $\alpha\rightarrow 0$ (the $In\ddot{o}n\ddot{u}$-$Wigner$ contraction [16]). Finally, by Eq. (A.2), the scalar product of the generators of the algebra (A.6) is given by
@@
or, in compact form  [^6] The numerical coefficient emerging in equation (A.7), can be reabsorbed by redefining the coupling constant of the action.
$(P_{a},P_{b})$ =±$=\pm|\alpha|\eta_{ab}=\alpha\eta_{ab},$ (A.8)

### v3

@@ 4.2 Recovering gravitational dynamics
We are now able to show that the Euler-Lagrange equations of the Yang-Mills action in (2.8), reduced to  (3.18), (3.19), are consistent with gravitational equations once the gauge fields are identified as described in (4.13). Equations (3.18) and(3.19), in components, read
@@
$\left(\frac{1}{2}\mathcal{D}_{\sigma}^{\varpi}\tilde{\mathfrak{T}}^{a}_{\rho\lambda}+\frac{1}{4}\mathfrak{R}_{[\mu\nu]}^{ab}\vartheta_{b\sigma}\varepsilon^{\mu\nu}_{\rho\lambda}+\frac{\lambda}{2}\vartheta_{b\sigma}\vartheta^{a}_{\mu}\vartheta^{b}_{\nu}\varepsilon^{\mu\nu}_{\rho\lambda}\right)\varepsilon^{\alpha\sigma\rho\lambda}$ $=\chi\Theta^{a}_{[\sigma\rho\lambda]}\varepsilon^{\alpha\sigma\rho\lambda},$ (4.15)
$\tilde{\mathfrak{T}}^{b}_{\rho\lambda}\vartheta^{a}_{\sigma}\varepsilon^{\alpha\sigma\rho\lambda}$ $=\chi\Sigma^{ab}_{[\sigma\rho\lambda]}\varepsilon^{\alpha\sigma\rho\lambda},$ (4.16)
@@
where we defined $\tilde{\mathfrak{T}}^{a}_{\rho\lambda}=\frac{1}{2}\varepsilon^{\mu\nu}_{\rho\lambda}\mathfrak{T}^{a}_{[\mu\nu]}$ and we used the explicit expressions $\mathfrak{R}^{ab}_{[\mu\nu]}=\partial_{\mu}\varpi^{ab}_{\nu}-\partial_{\nu}\varpi^{ab}_{\mu}+\varpi^{a}_{\mu\,c}\varpi^{cb}_{\nu}-\varpi^{a}_{\nu\,c}\varpi^{cb}_{\mu}$, $\mathfrak{T}^{a}_{[\mu\nu]}=\partial_{\mu}\vartheta^{a}_{\nu}-\partial_{\nu}\vartheta^{a}_{\mu}+\varpi^{a}_{\mu\,c}\vartheta^{c}_{\nu}-\varpi^{a}_{\nu\,c}\vartheta^{c}_{\mu}$. Now, using the definitions of the sources given in (3.6) and (3.7), in the limit $\alpha\to 0$ where the fields acquire a geometric interpretation, the Eqs. (4.15) and (4.16) can be written as
$\mathcal{D}_{\mu}^{\omega}\,T_{[\beta\nu]}^{a}\eta^{\mu\nu}+R^{ab}_{[\beta\nu]}\theta_{b\mu}\eta^{\mu\nu}+\lambda(\theta_{b\mu}\theta^{a}_{\beta}\theta^{b}_{\nu}-\theta_{b\mu}\theta^{a}_{\nu}\theta^{b}_{\beta})\eta^{\mu\nu}=\,$ $\chi\mathscr{T}^{a}_{\beta},$ (4.17)
$2\theta^{a}_{\mu}T^{b}_{[\beta\nu]}\eta^{\mu\nu}=\,$ $\chi\mathcal{S}^{ab}_{\beta}.$ (4.18)
In the same limit, the geometric field $\theta^{a}_{\mu}$ and its dual $\theta^{a\mu}=\eta^{\mu\nu}\theta^{a}_{\nu}$ can be used to define a new metric field
$g^{\mu\nu}=\eta_{ab}\theta^{a\mu}\theta^{b\nu}\,$, (4.19)
with inverse $g_{\mu\nu}=[g^{\mu\nu}]^{-1}$. Now the dual of $\theta^{a\mu}$ with respect to the new metric field is given by
$\overline{\theta}^{a}_{\mu}=g_{\mu\nu}\theta^{a\nu}$ (4.20)
with ${\theta}^{\mu}_{a},\overline{\theta}^{a}_{\mu}$, by construction, satisfying the tetrad relations
$\overline{\theta}^{a}_{\mu}\,\theta^{\nu}_{a}=\delta^{\nu}_{\mu}\,\,,\quad\overline{\theta}^{a}_{\mu}\,\theta^{\mu}_{b}=\delta^{a}_{b}.$ (4.21)
In terms of the new fields, Eqs. (4.17) and(4.18) take the form
@@
$R^{a}_{\beta}-\frac{1}{2}\overline{\theta}^{a}_{\beta}R+\Lambda^{\rho}_{\beta}\overline{\theta}^{a}_{\rho}=\chi(\mathscr{T}^{a}_{\beta}-\frac{1}{2}\overline{\theta}^{a}_{\beta}\mathscr{T})-\mathcal{D}_{\omega}^{\nu}T^{a}_{[\beta\nu]}+\frac{1}{2}\overline{\theta}^{a}_{\beta}(\theta^{\lambda}_{c}\mathcal{D}_{\omega}^{\sigma}T^{c}_{[\lambda\sigma]}),$ (4.22)
$T^{b}_{[\beta\nu]}=\frac{\chi}{2}\,\overline{\theta}_{a\nu}\mathcal{S}^{ab}_{\beta},$ (4.23)
with
$\Lambda^{\rho}_{\beta}=\lambda(gg^{\rho}_{\beta}-g^{\alpha}_{\beta}g^{\rho}_{\alpha})-\frac{1}{2}\lambda(g^{2}-g^{\alpha}_{\lambda}g^{\lambda}_{\alpha})\delta^{\rho}_{\beta},$ (4.24)
and we use $g=\eta_{\mu\nu}g^{\mu\nu}$ and $g^{\alpha}_{\mu}=\eta_{\mu\beta}g^{\beta\alpha}$. Moreover, $R^{a}_{\beta}=\theta^{\nu}_{b}R^{ab}_{[\beta\nu]}$ is the Ricci tensor, $R=\theta^{\beta}_{a}\theta^{\nu}_{b}R^{ab}_{[\beta\nu]}$ is the scalar curvature, while $\mathscr{T}=\theta^{\beta}_{a}\mathscr{T}^{a}_{\beta}$ is the trace of the energy-momentum tensor. This shows that, in the case where torsion is algebraically related to spin, the equations of motion for the curvature exhibit an additional source term in the energy-momentum tensor, which depends on the spin tensor. Moreover, the cosmological term becomes nontrivial, including contributions from self-interactions of the induced metric field and coupling terms with the Minkowski background metric.
@@ 5 Discussion
As for the parameters of the theory, in the limit $\alpha\rightarrow 0$ we have $\chi=-{2g^{2}}/{\hbar\lambda}={16\pi G_{N}}/{c^{3}}$. It follows that $\lambda$ and $g^{2}$ can be expressed in relation to one another through the squared Planck length $l^{2}_{p}=\hbar G_{N}/c^{3}$. As a consequence of the relation between $g^{2}$ and $\lambda$ and the convention adopted in the construction of the action, $\lambda$ must be negative, which in turn constrains $\alpha$ to be negative, as shown in Eq (2.2). This selection leads to $SO(2,3)$ as the family gauge structure group. However, adopting a different sign convention for the action would have led to a positive $\alpha$, thereby selecting the $SO(1,4)$ family. This highlights the fact that the sign of $\alpha$ does not play an essential role in the construction, which can be developed in a fully general manner for both the de Sitter and anti-de Sitter cases.
@@
A few more comments are in order. As discussed in [PhysRevLett.80.4851], the reliance of modern physics on two closely related but separate principles, i.e. general covariance and gauge invariance, may be unsatisfactory. Unifying these principles into a single framework is a task, and two main approaches emerge: one views gauge symmetry as a manifestation of general covariance in higher dimensions, while the other interprets general covariance as a form of gauge symmetry. The results presented in this paper go in the direction of the second approach. This occurs in the limit $\alpha\rightarrow 0$, where the gauge nature of the theory changes, providing a geometric interpretation of gauge potential components as the tetrad and the Lorentz connection.
@@
Since the model is obtained from a standard Yang-Mills action on Minkowski spacetime, well-established quantization techniques of non-Abelian gauge theories apply, although the issues related to the noncompactness of the structure group remain open, in terms of both unitarity and renormalizability of the quantum theory [Tseytlin:1981nu]. The results for the gravity sector of the theory should be analized while performing the limit $\alpha\rightarrow 0$.
@@
The possible physical meaning of the proposed mechanism, which reduces the gauge dynamics to the gravitational sector, deserves further investigation. One possibility could be a Higgs-like mechanism, which would involve a spontaneous breaking of the gauge symmetry to the residual Lorentz gauge symmetry, independently of the $\alpha$ limit (see e.g. [Ivanenko:1983fts, Percacci:1984bq]). Another possibility concerns the idea that the gauge parameters of the theory, $\lambda$ and $g$, be renormalized during quantization and would exhibit a running dependence on $\alpha$. The latter, as a dimensionless parameter, could be linked to the energy scale only if the Planck length is considered as a fundamental scale. A running of the gauge parameters $\lambda$ and $g$ could cause them to weigh the various tensors in the Euler-Lagrange equations differently, making some of them negligible compared to others. In other words, the dynamics we have described could emerge as effective dynamics resulting from the quantum nature of the theory. Finally, the meaning of the contraction parameter $\alpha$ and its interplay with the cosmological constant have to be understood, also in relation with the renormalization group flow.
As a final speculation, we observe that, since the dynamical metric is quadratic in the sector of P-translations of the gauge algebra, the undelying gauge theory could be a natural candidate for a double copy formulation of gravity [Campiglia:2021srh, Monteiro:2011pc, BCJ]. Focused work in this direction would require specifying solutions of the parent YM mapped to the gravitational sector. In this case, the interest of our group-theoretic reduction lies in the fact that the algebra characterizes the field solution whereas, in the double copy procedure, the gauge algebra of the YM field is generally unspecified. In this sense, recent work [LopesCardoso:2024ttc] investigates the relation between gravity and self-dual YM (SDYM) in 4D by dimensional reduction, via a shared description for specific sets of solutions of the two theories in terms of a 2D integrable Breitenlohner-Maison system and a single-copy lift back to 4D SDYM. In this case, the Breitenlohner-Maison procedure itself determines the Yang-Mills gauge group, encoding aspects of the gravitational solution in the Yang-Mills gauge group generators supporting the solution. We plan to address the possible connection with the double copy scenario and in particular with the findings of [LopesCardoso:2024ttc] in future work.
@@ Appendix A THE (A)dS GROUPS
In this appendix we fix the notation and our convention for the structure groups SO(p,q)SO({p},{q}) of the Yang-Mills model analyzed in the paper. These are the pseudo-orthogonal groups SO(p,q)SO({p},{q}) with $p+q=5$, namely the special Lie groups of linear transformations of a real five-dimensional vector space that leave invariant a nondegenerate symmetric bilinear form $\eta$ of signature $(p,q)$. The algebra $\mathfrak{so}(p,q)$ is provided by generators $M_{AB}$, $A,B=0,\dots,4$, and Lie brackets
$M_{AB},M_{CD}]=\eta_{AD}M_{BC}-\eta_{AC}M_{BD}+\eta_{BC}M_{AD}-\eta_{BD}M_{AC}:=f^{KL}_{AB,CD}M_{KL}$ (A.1)
@@
Specifically we are interested in the de Sitter group $SO(1,4)$ preserving the quadratic form $\eta_{AB}=diag(1,-1,-1,-1,-1)$ and the anti-de Sitter group $SO(2,3)$ preserving $\eta_{AB}=diag(1,-1,-1,-1,1)$. Our choice of basis for their Lie algebra is the one that singles out the Lorentz subalgebra. Indeed, if $a,b=0,1,2,3$ we can write
@@
where + and - refer, respectively, to the de Sitter and anti-de Sitter algebra. Hence we define
@@
where $\eta_{ab}$ is the Minkowski metric in 4D with signature (+,,,)(+,-,-,-) while $K^{cd}_{a,b}=\frac{1}{2}(\delta^{c}_{a}\delta^{d}_{b}-\delta^{c}_{b}\delta^{d}_{a})$. The parameter $\alpha$ can take positive, negative, or zero values. The Poincaré algebra $\mathfrak{iso}(1,3)$ can be derived from $\mathfrak{so}(1,4)$ or $\mathfrak{so}(2,3)$ in the limit of $\alpha\rightarrow 0$ (the $In\ddot{o}n\ddot{u}$-WignerWigner contraction [Inonu]). Finally, by Eq. (A.2), the scalar product of the generators of the algebra (A.6) is given by
@@
or, in compact form  [^6] The numerical coefficient emerging in Eq (A.7), can be reabsorbed by redefining the coupling constant of the action.
$(P_{a},P_{b})$ $=\pm|\alpha|\eta_{ab}=\alpha\eta_{ab},$ (A.8)